Test your knowledge:
molecule with 5000 base pairs (bp) - dsDNA
1. # nucleotides?
2. # of sugars?
3 # of helical turns? . # of helical turns?
4. Total length?
Food for thought: How many phosphodiester bonds?

Test your knowledge:
1. If a dsDNA molecule has 10% G how much
T is present?
Answer #1

1. Nucleotides are the building blocks of nucleic acids; they are composed of three subunit molecules: a nitrogenous base, a five-carbon sugar (ribose or deoxyribose), and at least one phosphate group.

As here 5000 base pairs are presesent that means total nucleotides will be 5000*2 =10000

2. Each nucleotide molecule contain a sugar molecule.

So, total number sugars are – 10000.

3. One turn requires 10 bp.

So, 5000 bp contain 5000/10 = 500 helical turn.

4. DNA helix has single turn 3.4nm in length.

So, 500 turns have [3.4*500] = 1700 nm length.

5. A dsDNA molecule have 10% G. So, C will also 10% as G=C.

Now total G+C = 20%

So, Total A+T = (100-20) = 80%

T is present – 80/2 = 40%